OpenStudy (anonymous):

what would be equivalent to f(x) = |x + 8| - 3 ?

OpenStudy (anonymous):

Very likely a function defined by pieces. f(x) = x + 8 - 3 = x+5 if |x+8| > 0 or f(x) = -(x+8) - 3 = -x -11 if |x+8| < 0 if I recall this correctly. Might want to double check if possible.

OpenStudy (anonymous):

Actually, it's only if x + 8 < 0, I misplaced the | | in both cases.

OpenStudy (anonymous):

@bmp these are the answer choices A. f(x) = x + 5 , if x > 3 -x - 11 , if x < 3 B. f(x) = x + 5 , if x > -8 -x - 11 , if x < -8 C. f(x) = -x - 5 , if x > -8 x + 11 , if x < -8 D. f(x) = -x - 5 , if x > 3 x + 11 , if x < 3

OpenStudy (anonymous):

B it's then, if you solve for x + 8 > 0 you'll get x > -8. Similarly, if you solve x + 8 < 0, you'll get x < -8.